Delos Debt Renegotiations​ (A). Delos borrowed euro€80 million two years ago. The loan​ agreement, an amortizing​ loan, was for six years at 8.622​% interest per annum. Delos has successfully completed two years of​ debt-service, but now wishes to renegotiate the terms of the loan with the lender to reduce its annual payments.

a. What were​ Delos's annual principal and interest payments under the original loan​ agreement?

b. After two years debt​ service, how much of the principal is still​ outstanding?

c. If the loan were restructured to extend another two​ years, what would be the annual paymentslong dash—principal

and​ interest? Is this a significant reduction from the original​ agreement's annual​ payments?
Answer #1

Delos borrowed amount= €80 million

Interest rate on loan = 8.622%

a). Calculating the annual payments of loan:

Annual Payment = P*r* (1+r) (1+r) - 1

where, P = Loan amount =  €80 million

r = periodic interest rate = 0.08622

n = no of years = 6

(1 +0.08622) Annual Payment = 80million * 0.08622 * - (1 + 0.08622)6 – 1

1.64250526 Annual Payment = 6897600 * 1,64250526-1

Annual Payments =  € 17,633,076.31

Now, Calculating the Delos's annual principal and Interest payment under the original loan​ agreement by providing the amortization table:

Year Opening bal. Payment Interest amount Principal Amount Closing Bal.
1     80,000,000.00    17,633,076.31            6,897,600.00                10,735,476.31       69,264,523.69
2     69,264,523.69    17,633,076.31            5,971,987.23                11,661,089.08       57,603,434.61
3     57,603,434.61    17,633,076.31            4,966,568.13                12,666,508.18       44,936,926.43
4     44,936,926.43    17,633,076.31            3,874,461.80                13,758,614.51       31,178,311.92
5     31,178,311.92    17,633,076.31            2,688,194.05                14,944,882.26       16,233,429.67
6     16,233,429.67    17,633,076.31            1,399,646.31                16,233,430.00                        (0.34)

b). After two years debt​ service, how much of the principal is still​ outstanding as per the above amortization table

=57,603,434.61

c). As the loan is restructured with amount € 57,603,434.61 for another additional 2 years making 6 years loan tenure (4 of original + 2 additional).

Calculating the annual payments of the loan:

Annual Payment = P*r* (1+r) (1+r) - 1

where, P = Loan amount =  € 57,603,434.61

r = periodic interest rate = 0.08622

n = no of years = 6

(1 + 0.08622) Annual Payment = 57, 603, 434.61 * 0.08622* (1 + 0.08622)6 – 1

1.64250526 Annual Payment = 4966, 568.13 * 1.64250526 - 1

Annual Payments = € 12,696,571.98

Now, Calculating the Delos's annual principal and Interest payment under the revised loan​ agreement by providing the amortization table:

Year Opening bal. Payment Interest amount Principal Amount Closing Bal.
1 57,603,434.61    12,696,571.98            4,966,568.13                  7,730,003.85       49,873,430.76
2     49,873,430.76    12,696,571.98            4,300,087.20                  8,396,484.78       41,476,945.98
3     41,476,945.98    12,696,571.98            3,576,142.28                  9,120,429.70       32,356,516.28
4     32,356,516.28    12,696,571.98            2,789,778.83                  9,906,793.15       22,449,723.14
5     22,449,723.14    12,696,571.98            1,935,615.13                10,760,956.85       11,688,766.29
6     11,688,766.29    12,696,571.98            1,007,805.43                11,688,766.55                        (0.26)

Interest amount = Opening balance*Monthly Interest rate

Principal amount = Payment - Interest amount

Closing balance = Opening Balance + Interest - Payment

- So, significant reduction from the original​ agreement's annual​ payments = € 17,633,076.31 - € 12,696,571.98

= € 4,936,504.33